OpenStudy (nbarrera):

PLEASE HELP ME! WILL MEDAL U! What number would you have to add to both sides to complete the square? x^2 – 16x = –8 A. 64 B. –16 C. 16 D. –64

OpenStudy (anonymous):

I believe it is -16

OpenStudy (narissa):

ask google

OpenStudy (nbarrera):

really?

OpenStudy (welshfella):

x^2 - 16x = (x -8) ^2 - k [its -8 because we take half of -16) what do you think the value of k is?

OpenStudy (anonymous):

16

OpenStudy (welshfella):

HInt: if we expand (x-8)^2 we get x^2 - 16x + 64

OpenStudy (nbarrera):

so 16?

OpenStudy (welshfella):

no not 16

OpenStudy (narissa):

@welshfella can i tag u and can u help me when ur finished here please

OpenStudy (welshfella):

ok

OpenStudy (welshfella):

x^2 - 16x + 64 + k = x^2 - 16 k = ?

OpenStudy (nbarrera):

-8

OpenStudy (welshfella):

no you need to get rid of the + 64 so as to make right side = x^2 - 16x so what do you add to 64 to make 0?

OpenStudy (welshfella):

check out my last post

OpenStudy (nbarrera):

-64

OpenStudy (welshfella):

yes! thats your answer -64

OpenStudy (welshfella):

x^2 - 16x + 64 - 64 = x^2 - 16

OpenStudy (welshfella):

x^2 - 16x = -8 (x -8)^2 - 64 = -8 -64 (x - 8^2 = -8 - 64 + 64 (x -)8^2 = -8